a current of 1.5A was passed through an electrolytic cell containing dilute tetraoxosulphateVIacid for 40 minutes . calculate the volume of gas at s.t.p .(F=96500C ,GMV=22.4)
1
Expert's answer
2019-09-03T02:33:34-0400

Total charge passed "=" "Q=It=1.5\times 40\times 60=" "3600C"

Now,

"m=\frac{MQ}{nF}"

Here,

"m=mass obtained"

"M=Molarmass"

Now, "No. of mole=" "\frac{m}{M}=\frac{Q}{nF}"

"F=96500C"


equation for Hydrogen:

"2H^+(aq)+2e^-\to" "H_2(g)"

"n=2"

Moles of hydrogen obtained"=\frac{Q}{nF}=\frac{3600}{2\times96500}=0.01865mole"

Volume of Hydrogen obtained"=0.01865\times22.4=0.418l" "("Molar Volume at S.T.P. "=22.4l" ")"

equation for Oxygen:

"4OH^-(aq)\to2H_2O(l)+4e^-"

"n=4"

Moles of Oxygen obtained"=\frac{Q}{nF}=" "\frac{3600}{4\times96500}=0.00933mole"

Volume of Oxygen obtained "=0.00933\times 22.4=0.209l"

Total volume obtained"=" Volume of Hydrogen "+" Volume of Oxygen "=0.418+0.209=0.627l"